OpenStudy (anonymous):

Answerr question for me? g(x) = 64(0.5x)

OpenStudy (mathstudent55):

What is the question?

OpenStudy (anonymous):

Is there any number for the variable x?

OpenStudy (anonymous):

@shybearr15 are u still there?

OpenStudy (anonymous):

Yes, sorry. I was busy. The whole question is: x f(x) = 2.5x – 10.5 ---------------------- g(x) = 64(0.5x) 2 2.5x – 10.5 = 64 x)" 3 (0.5x) 2.5x - 10.5 = 32x x)" 4 2.5x - 32x = 10.5 x)" 5 -29.5x = 10.5 x)" 6 x = -10.5/29.5 x)" Do you understand this? I do not. :/

OpenStudy (anonymous):

@Chocolate1013

OpenStudy (anonymous):

So what do you have to do? It doesn't even make sense...

OpenStudy (anonymous):

All it says is Complete the table below to solve the equation 2.5x – 10.5 = 64(0.5x).

OpenStudy (anonymous):

I have no idea. im confused too... :/

OpenStudy (anonymous):

Who?

OpenStudy (anonymous):

@amistre64 @SolomonZelman @phi

OpenStudy (anonymous):

Algebra is my worst subject....

OpenStudy (anonymous):

@Ashleyisakitty Can you help?